# A countable-support symmetric iteration separating $\mathsf{PP}$ from $\mathsf{AC}$

### v5

@@ 1 Introduction
Assume $V\models\mathsf{ZFC}$. In a class forcing extension $V[G]$ obtained by an $\mathrm{Ord}$-length countable-support symmetric iteration over a Cohen symmetric seed model $\mathcal{N}$, the final symmetric model $\mathcal{M}$ satisfies
@@
The iteration is an instance of the general countable-support symmetric-iteration framework developed in [5]; we use its successor and limit constructions, the $\omega_{1}$-completeness and normality facts for the generated filters, and the $\mathsf{ZF}/\mathsf{DC}$ preservation theorems. What is specific to the present manuscript is the package bookkeeping for $\mathsf{PP}^{\mathrm{split}}\!\restriction T$ and $\mathsf{AC}_{\mathsf{WO}}$ and the orbit-symmetrization required to obtain genuine automorphisms of the successor iterands (Remark 2.4).
@@
Section 2 fixes the metatheoretic conventions and records the countable-support symmetric-iteration API items imported from [5]. Section 3 constructs the seed model $\mathcal{N}$ and establishes $\mathsf{DC}$, $\neg\mathsf{AC}$, and $\mathsf{SVC}(S)$ at stage 0. Section 4 recalls Ryan–Smith localization and isolates the fixed parameter $T$ together with the localized principles. Subsections 4.1 and 4.2 define the package forcings used at successor stages. Section 5 defines the class-length symmetric iteration and verifies the final-stage propositions, culminating in Theorem 5.78.
@@ 2 Preliminaries and metatheoretic conventions
We follow the forcing, automorphism, symmetric-extension, and countable-support symmetric-iteration framework developed in [5], based on Karagila’s iteration scheme [9]. This section fixes the metatheory and notation used throughout, and records precisely which results are imported from [5].
@@ 2.1 Background theory and conventions
All forcing notions, names, automorphism groups, filters, and supports are constructed in a fixed background universe $V$.
  • For the set-forcing and symmetric-extension statements used at bounded stages, we work over $V\models\mathsf{ZF}$.
  • Whenever we invoke preservation of dependent choice for a bounded stage symmetric model, we assume the background satisfies $V\models\mathsf{ZFC}$. Throughout this paper, $\mathsf{DC}$ means $\mathsf{DC}_{\omega}$ (dependent choice for $\omega$-sequences, in the standard set formulation used in [5, §3.5]).
  • For the class-length recursion through $\mathrm{Ord}$, we assume a background theory adequate for class recursion (e.g. $V\models\mathsf{GBC}+\mathsf{ETR}$), and we assume Global Choice only when explicitly invoked for bookkeeping.
When we say “countable”, $\omega_{1}$-complete”, etc., this is always with respect to the background universe $V$.
@@ 2.2 Restrictions, supports, and stage generics
We follow the convention of [5, Remark 2.2] and do not overload restriction/projection notation.
@@
Let $G\subseteq\mathbb{P}_{\mathrm{Ord}}$ be $V$-generic for the class-length iteration forcing. For each $\alpha\leq\mathrm{Ord}$, write $G_{\alpha}:=G\cap\mathbb{P}_{\alpha}$ for the induced $V$-generic filter on the initial segment forcing.
2.3 Imported iteration results from [5]
@@
    Existence of set- and class-length iterations. The countable-support symmetric iteration recursion can be carried out through any set ordinal, and (under the background class-recursion assumptions used there) through $\mathrm{Ord}$; see [5, §Countable-support symmetric iterations and §Class-length iterations].
@@
    Limit filters. The limit-stage filter used in this manuscript is the modified filter $\tilde{\mathscr{F}}^{*}_{\lambda}$ (Definition 5.50). Its normality and $\omega_{1}$-completeness are proved here (Lemma 5.55). We only use [5] for the general iteration template and projection/coherence infrastructure leading to the limit stage.
@@
We stress that [5] does not assert $\mathsf{DC}$ for the final class-length symmetric model in general. In this paper, $\mathsf{DC}$ for the final model is obtained by combining item (d) at bounded stages with the stage-localization lemma for set-sized $\mathbb{P}_{\mathrm{Ord}}$-names (Lemma 5.67).
@@ 5 Iterating the packages over $\mathcal{N}$
    Black-box symmetric iteration. Subsection 5.5 invokes the symmetric-iteration framework to assemble the class-length iteration $\langle\mathbb{P}_{\alpha},\mathscr{G}_{\alpha},\tilde{\mathscr{F}}^{*}_{\alpha}:\alpha\in\mathrm{Ord}\rangle$, with $\mathbb{P}_{0}=\operatorname{Add}(\omega,\omega_{1})$ and successor stages $\mathbb{P}_{\alpha+1}=\mathbb{P}_{\alpha}*\dot{\mathbb{S}}_{\alpha}$ where $\dot{\mathbb{S}}_{\alpha}$ is the finite product of the scheduled orbit package factors. By Lemma 5.65, the canonical $\mathbb{P}_{\alpha}$-names for these orbit-package iterands are hereditarily symmetric, so the two-step presentation is well-defined in the stage ground $M_{\alpha}$.
@@
We now describe the symmetric iteration that forces the two conjuncts $\mathsf{PP}^{\mathrm{split}}\!\restriction T$ and $\mathsf{AC}_{\mathsf{WO}}$ over the Cohen symmetric seed model $\mathcal{N}$, while maintaining $\neg\mathsf{AC}$ (witnessed by the non-well-orderability of $A$). All technical iteration infrastructure is imported via the API of Remark 2.4 (in particular, existence of the class-length iteration is used exactly under the hypotheses of [5, Theorem 3.25]). The purpose of this section is to isolate the package systems, the bookkeeping targets, and the verification that the final symmetric model satisfies the desired principles.
@@ 5.1 Package symmetric systems at a stage
The countable completeness of the diagonal stage filters is what allows the CSI iteration theorem to handle limit stages. Concretely, the limit-step existence theorem for iterations of symmetric systems requires $\omega_{1}$-complete filters (CSI, [5, Theorem 3.13]), and we have imported this requirement as item (c) in Remark 2.4.
@@ 5.2 Iteration-level diagonal automorphisms
This is exactly the successor-step group structure supplied by the countable-support symmetric iteration framework (Remark 2.4, cf. [5, Def. 3.1]). Concretely:
@@
For $\gamma<\beta$ we write $\pi^{\gamma}_{\beta}:\mathscr{G}_{\beta}\to\mathscr{G}_{\gamma}$ for the canonical projection homomorphisms of the iteration groups (cf. [5, Def. 3.1]). These satisfy coherence: for all $\gamma<\beta<\lambda$,
@@
(i) This is immediate from Definition 5.23 together with the coherence of the projections (Remark 5.22); cf. also [5, Def. 3.1].
@@
At stage 0 there are no package coordinates. At a successor stage $\beta+1$, write $k=\mathrm{st}(h)\cdot u$ as in Lemma 5.21, where $h=\pi^{\beta}_{\beta+1}(k)\in\mathscr{G}_{\beta}$ and $u\in K_{\beta+1}$ acts trivially on the $\mathbb{P}_{\beta}$-part. By the inductive hypothesis, $h$ induces a bijection of $\mathrm{Coords}_{\beta}$. The standard lift $\mathrm{st}(h)$ extends this action and fixes every element of $\mathrm{NewCoords}_{\beta+1}$. On the other hand, $u$ fixes $\mathrm{Coords}_{\beta}$ pointwise and acts on the successor iterand $\dot{\mathbb{Q}}_{\beta}$ by compositions of package automorphisms (Lemma 5.8) and diagonal-cancellation automorphisms (Definition 5.11), each of which induces a bijection on the new coordinate index set. Therefore $k$ induces a bijection on $\mathrm{Coords}_{\beta+1}=\mathrm{Coords}_{\beta}\dot{\cup}\mathrm{NewCoords}_{\beta+1}$.
At a limit stage $\eta$, elements of $\mathscr{G}_{\eta}$ are coherent inverse-limit elements whose restrictions to earlier stages determine them uniquely (cf. [5, Def. 3.1]). The induced bijection on $\mathrm{Coords}_{\eta}=\bigcup_{\beta<\eta}\mathrm{Coords}_{\beta}$ is obtained by taking the union of the stagewise bijections, and is again a bijection. Consequently, if $D$ is countable then so is $k\cdot D$.
@@
Limit $\eta\leq\lambda$. Let $\eta$ be limit and assume $\langle k_{\beta}:\beta<\eta\rangle$ has been constructed coherently. By the limit-stage group definition ([5, Def. 3.1]), elements of $\mathscr{G}_{\eta}$ are exactly the coherent threads $\langle h_{\beta}\rangle_{\beta<\eta}$ with $\pi^{\gamma}_{\beta}(h_{\beta})=h_{\gamma}$ for $\gamma<\beta<\eta$. Since $\langle k_{\beta}\rangle_{\beta<\eta}$ is such a coherent thread by construction, it defines a unique element $k_{\eta}\in\mathscr{G}_{\eta}$ with $\pi^{\beta}_{\eta}(k_{\eta})=k_{\beta}$ for all $\beta<\eta$. Since $\widehat{\pi}^{\,D_{\eta}}$ has the same restrictions to each $\mathbb{P}_{\beta}$ for $\beta<\eta$, we have $k_{\eta}=\widehat{\pi}^{\,D_{\eta}}$.

### v6

@@ 1 Introduction
Assume $V\models\mathsf{ZFC}$. Assume $V\models\mathsf{ZFC}$. Assume $V\models\mathsf{ZFC}$, and let $G\subseteq\mathbb{P}_{\mathrm{Ord}}$ be $V$-generic in the metatheoretic sense of Remark 2.1 for the $\mathrm{Ord}$-length countable-support symmetric iteration over the Cohen symmetric seed model $\mathcal{N}$. Then the final symmetric model $\mathcal{M}$ satisfies
@@
The bounded-stage successor and limit constructions used here follow the countable-support symmetric-iteration scheme developed in [5]. From that companion paper we use the stagewise group/projection infrastructure and the bounded-stage preservation results, especially $\mathsf{ZF}$ and $\mathsf{DC}$ at set stages. What is specific to the present manuscript is the explicit $\mathrm{Ord}$-length recursion for the present package forcings, together with the diagonal-lift/diagonal-cancellation machinery used to define the limit filters and to make the bookkeeping work for $\mathsf{PP}^{\mathrm{split}}\!\restriction T$ and $\mathsf{AC}_{\mathsf{WO}}$ (Remark 2.4).
@@
Section 2 fixes the metatheoretic conventions, records the bounded-stage iteration facts imported from [5], and explains how the present manuscript carries out the specific $\mathrm{Ord}$-length recursion. Section 3 constructs the seed model $\mathcal{N}$ and establishes $\mathsf{DC}$, $\neg\mathsf{AC}$, and $\mathsf{SVC}(S)$ at stage 0. Section 4 recalls Ryan–Smith localization and isolates the fixed parameter $T$ together with the localized principles. Subsections 4.1 and 4.2 define the package forcings used at successor stages. Section 5 defines the class-length symmetric iteration and verifies the final-stage propositions, culminating in Theorem 5.85.
@@ 2 Preliminaries and metatheoretic conventions
We follow the forcing, automorphism, symmetric-extension, and bounded-stage countable-support symmetric-iteration framework developed in [5], based on Karagila’s iteration scheme [9]. This section fixes the metatheory and notation used throughout, records precisely which bounded-stage results are imported from [5], and explains how the present paper carries out the specific $\mathrm{Ord}$-length recursion needed for the package iteration.
@@ 2.1 Background theory and conventions
All forcing notions, names, automorphism groups, filters, and supports are constructed over a fixed background set universe $V$.
  • For the set-forcing and symmetric-extension statements used at bounded stages, it suffices to work over $V\models\mathsf{ZF}$.
  • Whenever we invoke preservation of dependent choice for a bounded stage symmetric model, we assume $V\models\mathsf{ZFC}$. Throughout this paper, $\mathsf{DC}$ means $\mathsf{DC}_{\omega}$ (dependent choice for $\omega$-sequences, in the standard set formulation used in [5, Definition (Dependent Choice)]).
  • For the class-length recursion through $\mathrm{Ord}$ (bookkeeping and the definition of the iteration itself), we work in a background two-sorted structure $(V,\mathcal{C})$ satisfying a class theory adequate for transfinite recursion along $\mathrm{Ord}$ (e.g. $(V,\mathcal{C})\models\mathsf{GBC}+\mathsf{ETR}_{\mathrm{Ord}}$, or simply $\mathsf{GBC}+\mathsf{ETR}$ in the standard sense). When convenient we fix a class well-order of $V$ (equivalently, assume Global Choice in the background), but we use such Global Choice only for bookkeeping/canonical enumerations and explicitly indicate when it is invoked.
When we speak of a “$V$-generic” $G\subseteq\mathbb{P}_{\mathrm{Ord}}$, this is understood in the metatheory (equivalently, in an outer universe) as a subclass meeting every dense class $D\in\mathcal{C}$. We emphasize that we do not rely on a global forcing theorem for $\mathbb{P}_{\mathrm{Ord}}$: valuations of set-sized $\mathbb{P}_{\mathrm{Ord}}$-names localize to bounded forcing stages (Lemma 5.72), and set-sized final hereditarily symmetric names descend to bounded-stage hereditarily symmetric names (Lemma 5.73). Accordingly, set-theoretic arguments are carried out inside initial-segment extensions $V[G_{\alpha}]$ and their bounded-stage symmetric submodels.
Finally, when we say “countable”, $\omega_{1}$-complete”, etc., this is always with respect to the background universe $V$.
@@ 2.2 Restrictions, supports, and stage generics
We follow the notation discipline for forcing versus group restrictions from [5, Remark (Notation discipline for later sections)] and do not overload restriction/projection notation.
@@
Fix (in the metatheory) a $V$-generic subclass $G\subseteq\mathbb{P}_{\mathrm{Ord}}$ for the class-length iteration forcing. For each set ordinal $\alpha$, write $G_{\alpha}:=G\cap\mathbb{P}_{\alpha}$ for the induced $V$-generic filter on the initial-segment forcing.
2.3 Imported bounded-stage iteration results from [5]
@@
    Bounded-stage recursion and coherence. For each set ordinal $\lambda$, the countable-support symmetric iteration framework provides the bounded-stage forcing $\mathbb{P}_{\lambda}$, the group $\mathscr{G}_{\lambda}$, the stage filter, and the projection/coherence maps used throughout the set-stage analysis; see [5, §Limit Filters for Countable-Support Symmetric Iterations]. The present paper does not import a general class-length existence theorem from [5]; instead, the $\mathrm{Ord}$-length iteration used here is defined explicitly by recursion in the background metatheory described in Remark 2.1.
@@
    Limit filters. The limit-stage filter used in this manuscript is the modified filter $\tilde{\mathscr{F}}^{*}_{\lambda}$ (Definition 5.53). Its normality and $\omega_{1}$-completeness are proved here (Lemma 5.58). We only use [5] here for the bounded-stage projection/coherence infrastructure and the underlying countable-support limit construction.
@@
We stress that [5] is used here only at bounded stages. In particular, this paper does not import any general theorem asserting $\mathsf{DC}$ or $\mathsf{ZF}$ for the final class-length symmetric model. The final-model arguments are proved directly in the present manuscript: $\mathsf{DC}$ is obtained by combining item (d) with the stage-descent lemma for set-sized final hereditarily symmetric names (Lemma 5.73), and $\mathsf{ZF}$ for the final model is verified by the standard symmetric-extension argument applied to the final class system (Theorem 5.64 and Remark 5.63).
@@ 5 Iterating the packages over $\mathcal{N}$
    The $\mathrm{Ord}$-length recursion. Subsection 5.5 defines the specific $\mathrm{Ord}$-length iteration $\langle\mathbb{P}_{\alpha},\mathscr{G}_{\alpha},\tilde{\mathscr{F}}^{*}_{\alpha}:\alpha\in\mathrm{Ord}\rangle$ used in this paper, starting from $\mathbb{P}_{0}=\operatorname{Add}(\omega,\omega_{1})$ and setting $\mathbb{P}_{\alpha+1}=\mathbb{P}_{\alpha}*\dot{\mathbb{S}}_{\alpha}$ at successor stages, where $\dot{\mathbb{S}}_{\alpha}$ is the finite product of the scheduled orbit-package factors. The bounded-stage successor/limit infrastructure comes from Remark 2.4, while the transfinite recursion through $\mathrm{Ord}$ is carried out explicitly in the background metatheory from Remark 2.1. By Lemma 5.70, the canonical $\mathbb{P}_{\alpha}$-names for these orbit-package iterands are hereditarily symmetric, so the two-step presentation is well-defined in the stage ground $M_{\alpha}$.
@@
We now describe the symmetric iteration that forces the two conjuncts $\mathsf{PP}^{\mathrm{split}}\!\restriction T$ and $\mathsf{AC}_{\mathsf{WO}}$ over the Cohen symmetric seed model $\mathcal{N}$, while maintaining $\neg\mathsf{AC}$ (witnessed by the non-well-orderability of $A$). All technical bounded-stage iteration infrastructure is imported via the API of Remark 2.4. The $\mathrm{Ord}$-length iteration used in this section is then defined explicitly in the present manuscript by recursion in the background metatheory from Remark 2.1; it is not an instance of a general class-length existence theorem imported from [5]. The purpose of this section is to isolate the package systems, the bookkeeping targets, and the verification that the final symmetric model satisfies the desired principles.
@@ 5.1 Package symmetric systems at a stage
The countable completeness of the diagonal stage filters is what makes the bounded-stage countable-support limit construction compatible with the hereditary-symmetry and $\mathsf{DC}$-preservation arguments. In [5], $\omega_{1}$-completeness is a key property of the stage filters at limits (see the limit-filter construction and the theorem on limit filter properties), and we import that requirement here as item (c) in Remark 2.4.
@@ 5.2 Iteration-level diagonal automorphisms
This is exactly the successor-step group structure supplied by the bounded-stage countable-support symmetric iteration framework (Remark 2.4; cf. [5, §Framework and Setup]). Concretely:
@@
For $\gamma<\beta$ we write $\pi^{\gamma}_{\beta}:\mathscr{G}_{\beta}\to\mathscr{G}_{\gamma}$ for the canonical projection homomorphisms of the iteration groups (cf. [5, Definition (Countable-support iteration setup)]). These satisfy coherence: for all $\gamma<\beta<\lambda$,
@@
(i) This is immediate from Definition 5.23 together with the coherence of the projections (Remark 5.22); cf. also [5, Definition (Countable-support iteration setup)].
@@
At stage 0 there are no package coordinates. At a successor stage $\beta+1$, write $k=\mathrm{spl}(h)\cdot u$ as in Lemma 5.21, where $h=\pi^{\beta}_{\beta+1}(k)\in\mathscr{G}_{\beta}$ and $u\in K_{\beta+1}$ acts trivially on the $\mathbb{P}_{\beta}$-part. By the inductive hypothesis, $h$ induces a bijection of $\mathrm{Coords}_{\beta}$. The standard lift $\mathrm{spl}(h)$ extends this action and fixes every element of $\mathrm{NewCoords}_{\beta+1}$. On the other hand, $u$ fixes $\mathrm{Coords}_{\beta}$ pointwise and acts on the successor iterand $\dot{\mathbb{Q}}_{\beta}$ by compositions of package automorphisms (Lemma 5.8) and diagonal-cancellation automorphisms (Definition 5.11), each of which induces a bijection on the new coordinate index set. Therefore $k$ induces a bijection on $\mathrm{Coords}_{\beta+1}=\mathrm{Coords}_{\beta}\dot{\cup}\mathrm{NewCoords}_{\beta+1}$.
At a limit stage $\eta$, elements of $\mathscr{G}_{\eta}$ are coherent inverse-limit elements whose restrictions to earlier stages determine them uniquely (cf. [5, Definition (Countable-support iteration setup)]). The induced bijection on $\mathrm{Coords}_{\eta}=\bigcup_{\beta<\eta}\mathrm{Coords}_{\beta}$ is obtained by taking the union of the stagewise bijections, and is again a bijection. Consequently, if $D$ is countable then so is $k\cdot D$.
@@
Lemma 5.45 (Tail extension by identity).
Let $\beta\leq\lambda$ be stages of the iteration, where $\lambda$ may also be $\mathrm{Ord}$ in the background class metatheory. For every $k\in\mathscr{G}_{\beta}$, there exists $\hat{k}\in\mathscr{G}_{\lambda}$ such that:
  1. (i)
    $\pi^{\beta}_{\lambda}(\hat{k})=k$;
  2. (ii)
[…]
    $\hat{k}_{\gamma+1}=\mathrm{spl}(\hat{k}_{\gamma});$
  3. (iii)
    equivalently, $\hat{k}$ acts as $k$ on the $\mathbb{P}_{\beta}$-part and trivially on every iterand added at stages in $\beta,\lambda)$.
Proof.
If $\beta=\lambda$, take $\hat{k}:=k$. Assume $\beta<\lambda$.
We build by transfinite recursion a coherent thread $\langle k_{\gamma}:\beta\leq\gamma\leq\lambda\rangle$ such that $k_{\beta}=k$, each $k_{\gamma}\in\mathscr{G}_{\gamma}$, and $\pi^{\delta}_{\gamma}(k_{\gamma})=k_{\delta}$ whenever $\beta\leq\delta<\gamma\leq\lambda$.
Base stage. Set $k_{\beta}:=k$.
Successor stage $\gamma+1$. Given $k_{\gamma}\in\mathscr{G}_{\gamma}$, define
$k_{\gamma+1}:=\mathrm{spl}(k_{\gamma})\in\mathscr{G}_{\gamma+1},$
where $\mathrm{spl}$ is the splitting section from Lemma 5.21. Then
$\pi^{\gamma}_{\gamma+1}(k_{\gamma+1})=\pi^{\gamma}_{\gamma+1}(\mathrm{spl}(k_{\gamma}))=k_{\gamma},$
so coherence is preserved. By construction, $k_{\gamma+1}$ acts as $k_{\gamma}$ on the $\mathbb{P}_{\gamma}$-part and trivially on the new iterand at stage $\gamma$.
Limit stage $\eta\leq\lambda$. Assume $\langle k_{\gamma}:\beta\leq\gamma<\eta\rangle$ has been constructed coherently. By the limit-stage group definition used in the iteration, coherent threads determine unique elements of $\mathscr{G}_{\eta}$. Thus there is a unique $k_{\eta}\in\mathscr{G}_{\eta}$ such that
$\pi^{\gamma}_{\eta}(k_{\eta})=k_{\gamma}\qquad(\beta\leq\gamma<\eta).$
If $\eta=\mathrm{Ord}$, this is carried out in the same background metatheory used for the class-length bookkeeping (Remark 5.65).
Now set $\hat{k}:=k_{\lambda}$. Then (i) holds by construction. Clause (ii) is exactly the successor clause of the recursion. Clause (iii) is the equivalent operational description: at each successor step we extend by the splitting section, which acts trivially on the new iterand and preserves the previously constructed action. ∎
Corollary 5.46 (Identity extension for $\Delta^{\uparrow}$-groups).
Let $\beta\leq\lambda$, let $E\subseteq\omega_{1}$, and let $D\subseteq\mathrm{Coords}_{\beta}$. If $k\in\Delta_{\beta}^{\uparrow}(E,D)$, then there exists $\hat{k}\in\Delta_{\lambda}^{\uparrow}(E,D)$ such that
$\pi^{\beta}_{\lambda}(\hat{k})=k.$
Proof.
Let $\hat{k}\in\mathscr{G}_{\lambda}$ be the identity-tail extension of $k$ given by Lemma 5.45.
For the stage-0 part, by Definition 5.23 and projection coherence (Remark 5.22),
$\pi^{0}_{\lambda}(\hat{k})=\pi^{0}_{\beta}(\pi^{\beta}_{\lambda}(\hat{k}))=\pi^{0}_{\beta}(k)\in\operatorname{Fix}(E).$
For the coordinate part, let $d\in D\subseteq\mathrm{Coords}_{\beta}$. At stage $\beta$, the element $k$ fixes $d$ pointwise. At every successor step above $\beta$, the extension is obtained by the splitting section, which preserves the action on all previously existing coordinates and acts trivially on the new iterand. By induction on stages from $\beta$ up to $\lambda$, it follows that $\hat{k}$ fixes $d$ pointwise as well.
Hence $\hat{k}\in\Delta_{\lambda}^{\uparrow}(E,D)$ by Definition 5.29(A). ∎
Lemma 5.47 (Action on bounded names factors through projection).
Let $\beta\leq\lambda$ be stages, let $\tau$ be a $\mathbb{P}_{\beta}$-name, and let $k\in\mathscr{G}_{\lambda}$. Viewing $\tau$ as a $\mathbb{P}_{\lambda}$-name by canonical inclusion, one has
$k\cdot\tau\;=\;\pi^{\beta}_{\lambda}(k)\cdot\tau.$
Proof.
We argue by induction on the rank of $\tau$. Let $\langle\sigma,p\rangle\in\tau$. Since $\tau$ is a $\mathbb{P}_{\beta}$-name, we have $p\in\mathbb{P}_{\beta}$. The projection $\pi^{\beta}_{\lambda}(k)$ is, by definition, the restriction of $k$ to the $\mathbb{P}_{\beta}$-part, so
$k\cdot p\;=\;\pi^{\beta}_{\lambda}(k)\cdot p.$
By the induction hypothesis,
$k\cdot\sigma\;=\;\pi^{\beta}_{\lambda}(k)\cdot\sigma.$
Therefore the image of each pair $\langle\sigma,p\rangle$ under $k$ is exactly the same as its image under $\pi^{\beta}_{\lambda}(k)$, and hence $k\cdot\tau=\pi^{\beta}_{\lambda}(k)\cdot\tau$. ∎
@@
Limit $\eta\leq\lambda$. Let $\eta$ be limit and assume $\langle k_{\beta}:\beta<\eta\rangle$ has been constructed coherently. By the limit-stage group clause in the countable-support iteration setup ([5, Definition (Countable-support iteration setup)]), elements of $\mathscr{G}_{\eta}$ are exactly the coherent threads $\langle h_{\beta}\rangle_{\beta<\eta}$ with $\pi^{\gamma}_{\beta}(h_{\beta})=h_{\gamma}$ for $\gamma<\beta<\eta$. Since $\langle k_{\beta}\rangle_{\beta<\eta}$ is such a coherent thread by construction, it defines a unique element $k_{\eta}\in\mathscr{G}_{\eta}$ with $\pi^{\beta}_{\eta}(k_{\eta})=k_{\beta}$ for all $\beta<\eta$. Since $\widehat{\pi}^{\,D_{\eta}}$ has the same restrictions to each $\mathbb{P}_{\beta}$ for $\beta<\eta$, we have $k_{\eta}=\widehat{\pi}^{\,D_{\eta}}$.